OpenStudy (montesanna99):

The lengths of the sides of a triangle are in the ratio 5:6:7. Describe the length of the longest side if the perimeter is less than 54 cm

OpenStudy (aum):

The longest side will be \(\Large \frac{7}{(5+6+7)} * \text{ Perimeter }\) If \(\Large \text{ Perimeter } \lt 54~~ \text{cm}\)., then the longest side will have to be \(\Large \lt \frac{7}{(5+6+7)} * 54 = ?\)

OpenStudy (montesanna99):

21

OpenStudy (montesanna99):

How did u get the seven as the numerator

OpenStudy (aum):

The ratio of the sides of the triangle are: 5:6:7. The biggest of these numbers is 7 and therefore must represent the longest side. If we add up the three numbers we get 18. So the longest side must be 7/18 of the perimeter. But since the perimeter is less than 54, the longest side must be less than 7/18 * 54 = 21. So the longest side must be less than 21 cm.